OpenStudy (anonymous):

Simplify....1/y+1/x divided by 1/y-1/x

myininaya (myininaya):

\[\frac{\frac{1}{y}+\frac{1}{x}}{\frac{1}{y}-\frac{1}{x}}\]

myininaya (myininaya):

try multiplying top and bottom by xy to get rid of the compound fraction look

OpenStudy (anonymous):

so you should get an answer of (x+y)/(x-y)